what is the integral of ((e^x)/(1+x^2))-((2x*e^x)/((1+x^2)^2)) ?

Grade:11

1 Answers

Arun
6 years ago
Dear student
we know
integral of 
\tiny \intex [f(x) + f ’(x)] dx = ex f(x) + C
same case is here
if you look closely
(e^x)[1/(1+x^2) – 2x/(1+x^2)^2]\tiny \int = 
 =  (e^x)[1/(1+x^2) + C
Reagrds
Arun (askIITians forum expert)